OpenStudy (anonymous):

Solve -4p - 10(2p-3)= 22-7(1-2p). Please show all of your work.

OpenStudy (anonymous):

-4p-20p+30=22-7+14p -24p+30=15+14p -24p-14p=15-30 -38p=-15 38p=15 p=15/38

OpenStudy (anonymous):

thanks